# Scenario analysis with multivariate Bayesian machine learning models

### v1

@@
We present an econometric framework that adapts tools for scenario analysis, such as conditional forecasts and generalized impulse response functions, for use with dynamic nonparametric multivariate models. We demonstrate the utility of this approach through an exercise with simulated data, and three real-world applications: (i) scenario-based conditional forecasts aligned with Federal Reserve Bank stress test assumptions, (ii) measuring macroeconomic risk under varying financial conditions, and (iii) the asymmetric effects of US-based financial shocks and their international spillovers. Our results indicate the importance of nonlinearities and asymmetries in the dynamic relationship between macroeconomic and financial variables.
@@
Keywords: conditional forecast, generalized impulse response function, Bayesian additive regression trees, nonlinearities, (semi)-structural inference
Contact: Michael Pfarrhofer (michael.pfarrhofer@wu.ac.at), Department of Economics, WU Vienna University of Economics and Business. We would like to thank Niko Hauzenberger for useful comments. Disclaimer: The views expressed in this paper do not necessarily reflect those of the Oesterreichische Nationalbank or the Eurosystem.
@@ 1. Introduction
In this paper, we discuss how to conduct scenario analysis with multivariate time series models in a macroeconomic context when the functional form of the conditional mean is nonlinear and/or unknown. This situation can arise when relying on “traditional” nonlinear frameworks (e.g., in variants of threshold, regime-switching, or time-varying parameter models, see Fischer et al. (2023) for a recent example), but is virtually always the case in recently developed models that introduce Bayesian machine learning (ML) techniques to multivariate macroeconometric modeling (see, e.g., Huber and Rossini, 2022; Clark et al., 2023; Huber et al., 2023; Hauzenberger et al., 2024c).
@@
We use the term scenario analysis broadly to refer to different types of counterfactual experiments, including a version of conditional forecasts and variants of generalized impulse response functions (GIRFs, see Koop et al., 1996). While some aspects and issues in this context have been discussed in isolation in the aforementioned papers, there is no unified explicit treatment or comprehensive framework available yet. Inspired by the closely related work of Crump et al. (2025), which focuses on a linear Bayesian vector autoregression (VAR), this is the gap in the literature that we aim to fill with our paper.
@@
Conditional forecasts — the simulation of the future path of selected variables conditional on predefined scenarios encoded in constraints placed on observed variables — have been used by academics and practitioners since the 1980s in a linear VAR context (see, e.g., Doan et al., 1984), with subsequent refinements and improvements in computational efficiency (see, e.g., Waggoner and Zha, 1999; Jarociński, 2010; Bańbura et al., 2015). A recent and highly efficient computational approach is developed in Chan et al. (2025). Their framework exploits the properties of conditionally Gaussian errors in conjunction with the common assumption of a known linear conditional mean relationship. The full future path of restricted and unrestricted forecasts across multiple horizons can be jointly sampled from a high-dimensional (potentially truncated) multivariate Gaussian distribution.
@@
Breaking the assumption of linearity complicates matters. In particular, the presence of nonlinearities makes it difficult to derive multi-step ahead predictive distributions, and in many cases it will be outright impossible to obtain closed-form solutions. In this case, one can resort to recursive predictive simulation. This paper discusses how common approaches to conditional forecasting can be adapted to the context of nonparametric multivariate models. The proposed approach is quite general in the sense that it can be used in conjunction with most popular frameworks that introduce nonlinearities in the conditional mean of multivariate models. The key assumption is that the reduced form errors are conditionally multivariate Gaussian.
@@
Since the impulse response function (IRF), another key macroeconomic estimand, is commonly defined as the difference between forecasts conditioned on the values of different shocks, our proposed framework is also useful for semi-structural scenario analysis in a nonlinear context (see Antolin-Diaz et al., 2021, for an excellent discussion when using linear structural VARs, SVARs). Specifically, we revisit the issue of how to obtain dynamic causal effects, in the form of GIRFs, in a nonlinear and possibly nonparametric context. We examine both unrestricted GIRFs to shocks identified with approaches typically used in the linear SVAR literature, and restricted GIRFs, which can be used to investigate and quantify the contributions of specific sets of transmission channels in the propagation of structural shocks. The latter are obtained by partially matching the moments of the unconditional and conditional predictive distributions used to compute the GIRFs.
@@
We mentioned above that the proposed approach is quite generally applicable in multivariate models whenever one is willing to assume conditionally Gaussian innovations. This means that there is a large number of potential candidates for the respective conditional mean functions that can be assumed. These choices include (but are not limited to) regression tree-based implementations, Gaussian process priors, or neural networks (see Marcellino and Pfarrhofer, 2024, for a recent Bayesian overview). For our empirical work, we use Bayesian Additive Regression Trees (BART, Chipman et al., 2010) as a specific nonparametric implementation to infer functional relationships in a multivariate time series model. We pick this sum-of-trees model because tree-based approaches have proven particularly capable of producing accurate forecasts when used with time series data for the US economy (see, e.g., Medeiros et al., 2021; Goulet Coulombe et al., 2022; Clark et al., 2023; Goulet Coulombe, 2024), with datasets typically structured similarly to the one we use in this paper.
@@
Our approach to scenario analysis is developed under the assumption of a multivariate Gaussian error term with a general time-varying covariance matrix. From a practical and implementation perspective, we rely on a somewhat simplified version to capture time variation of the respective volatilities. The framework allows for flexible equation-by-equation estimation of the multivariate system and offers significant improvements in computational efficiency also in the context of conditional forecasts and GIRFs. In our empirical work, we capture heteroskedastic features of the data with a specification related to common volatility approaches (see Carriero et al., 2016), which also reflects recent tools used to address outliers during the Covid-19 pandemic (see, e.g., Carriero et al., 2024). It is worth noting that all methods also work with more sophisticated volatility models (see, e.g., Chan, 2023), albeit at the cost of increased computational burden.
@@
After investigating the performance our approach with synthetic data, we apply our framework in several related yet distinct cases. Our dataset comprises about $25$ quarterly macroeconomic and financial variables for the US economy ranging from the mid-$1970$s to the last quarter of $2023$. In one of our explorations, we also add some international variables and include data from the euro area (EA), Japan, and the United Kingdom (UK) in our model. The applied work assesses and illustrates the role of nonlinearities when interest centers on conditional forecasts, and we also explore asymmetries in the propagation of structural shocks of different signs and magnitudes. Specifically, we provide three empirical applications. First, inspired by Chan et al. (2025), we use a subset of the assumptions underlying the annual stress test conducted by the Federal Reserve System and compute conditional forecasts using soft constraints for different scenarios, comparing predictive densities from linear and nonlinear models. Second, reflecting the growth-at-risk literature (see Adrian et al., 2019), we study the counterfactual implications of varying financial conditions (imposing hard constraints over a period) on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock (as in, e.g., Barnichon et al., 2022), and compute GIRFs to shocks of different signs and magnitudes that are allowed to propagate internationally. We then use a restricted GIRF approach to gauge the role of spillovers and spillbacks.
@@
The rest of this paper is structured as follows. Section 2 lays out the baseline econometric framework and discusses an implementation using BART. Section 3 provides remarks about obtaining predictive inference with Monte Carlo methods. We discuss how to impose hard and soft constraints on forecasts, and how these constraints may be used to construct semi-structural scenarios through generalized impulse response functions. Section 4 provides an empirical illustration of the proposed methods. The last section concludes.
2. Econometric framework
@@
Let $\bm{y}_{t}=(y_{1t},\ldots,y_{nt})^{\prime}$ collect $n$ variables for $t=1,\ldots,T,$ and $\bm{x}_{t}=(\bm{y}_{t-1}^{\prime},\ldots,\bm{y}_{t-p}^{\prime})^{\prime}$ is a $k\times 1$ vector that stacks lags of $\bm{y}_{t}$, thus $k=np$. We use a multivariate model of the form:

### v2

@@
We present an econometric framework that adapts tools for scenario analysis, such as variants of conditional forecasts and impulse response functions, for use with dynamic nonparametric multivariate models. We demonstrate the utility of our approach with simulated data and three real-world applications: (1) scenario-based conditional forecasts aligned with Federal Reserve stress test assumptions, measuring (2) macroeconomic risk under varying financial conditions, and (3) asymmetric effects of US-based financial shocks and their international spillovers. Our results indicate the importance of nonlinearities and asymmetries in dynamic relationships between macroeconomic and financial variables.
@@
Keywords: conditional forecast, generalized impulse response function, Bayesian additive regression trees, nonlinearities, structural inference
Contact: michael.pfarrhofer@wu.ac.at, Department of Economics, WU Vienna University of Economics and Business. We would like to thank Niko Hauzenberger for useful comments. Disclaimer: The views expressed in this paper do not necessarily reflect those of the Oesterreichische Nationalbank or the Eurosystem.
@@ 1. Introduction
In this paper, we discuss how to conduct scenario analysis with multivariate time series models in a macroeconomic context when the functional form of the conditional mean is nonlinear and/or unknown. This situation can arise when relying on “traditional” nonlinear frameworks (e.g., in variants of threshold, regime-switching, or time-varying parameter models, see Fischer et al. (2023) for a recent example), but is virtually always the case in recently developed models that introduce Bayesian machine learning (ML) techniques to multivariate macroeconometric modeling (see, e.g., Huber and Rossini, 2022; Clark et al., 2023; Huber et al., 2023; Hauzenberger et al., 2024c). Our framework combines elements of structural vector autoregressions (SVARs) with nonlinear reduced form models, which is why we sometimes refer to aspects of our approach as “semi-structural.”
@@
We use the term scenario analysis rather broadly to refer to different counterfactual experiments, including a version of conditional forecasts (CFs) and variants of nonlinear impulse response functions (IRFs). While some aspects and issues in the Bayesian ML context have been discussed in isolation in the aforementioned papers, there is no unified explicit treatment or comprehensive framework available yet. Inspired by the related work of Crump et al. (2025), which focuses on a linear Bayesian vector autoregression (VAR), this is the gap in the literature that we aim to fill with our paper.
@@
CFs — simulating the future path of a subset of variables conditional on predefined scenarios encoded in constraints placed on other observed variables — have been used by academics and practitioners since the 1980s in a linear VAR context (see, e.g., Doan et al., 1984), with subsequent refinements and improvements in computational efficiency (see, e.g., Waggoner and Zha, 1999; Jarociński, 2010; Bańbura et al., 2015). A recent and highly efficient computational approach is developed in Chan et al. (2025). Their framework exploits the properties of conditionally Gaussian errors in conjunction with the common assumption of a known linear conditional mean relationship.
@@
Breaking the assumption of linearity complicates matters. In particular, there is no general causal representation of nonlinear time series as functions of structural shocks (see, e.g., Potter, 2000, for a discussion), and the presence of unknown nonlinearities makes it difficult to derive multi-step ahead predictive distributions. In many cases it is outright impossible to obtain closed-form solutions. In this case, one can resort to recursive predictive simulations. Indeed, this paper discusses how common approaches to (conditional) forecasting can be adapted to the context of nonparametric multivariate models for various types of scenario analyses based on Monte Carlo methods. The proposed approach is quite general in the sense that it can be used in conjunction with most popular frameworks that introduce nonlinearities in the conditional mean of multivariate models. The key assumption is that the errors are conditionally multivariate Gaussian.
@@
Since the IRF, another key macroeconomic estimand, can be defined as the difference between forecasts conditioned on the values of different shocks (for seminal work, see Gallant et al., 1993; Koop et al., 1996), our proposed framework is also useful for nonlinear semi-structural scenario analysis. For excellent discussions in the context of linear SVARs and partially nonlinear structural models, see Antolin-Diaz et al. (2021) and Gonçalves et al. (2021). Specifically, we revisit the issue of how to obtain dynamic causal effects, in the form of generalized IRFs (GIRFs), in models with nonlinearities of unknown form. Compared with IRFs from standard linear VARs, which are symmetric, shape invariant and history independent, our GIRFs do not feature these potentially restrictive properties. We examine both unrestricted GIRFs to shocks identified with approaches typically used in the linear SVAR literature, and restricted GIRFs, which can be used to investigate and quantify the contributions of specific sets of transmission channels in the propagation of shocks. The latter are obtained by partially matching the moments of the unconditional and conditional predictive distributions used to compute the GIRFs.
@@
We mentioned above that the proposed approach is quite generally applicable in multivariate dynamic models. This means that there is a large number of potential candidates for the respective conditional mean functions that can be assumed. These choices include (but are not limited to) regression tree-based implementations or Gaussian process priors (see Marcellino and Pfarrhofer, 2024, for a recent review). For our empirical work, we use Bayesian Additive Regression Trees (BART, Chipman et al., 2010) as a specific nonparametric implementation to infer functional relationships in a multivariate time series model. We pick this sum-of-trees model because tree-based approaches have proven particularly capable of producing accurate forecasts when used with time series data for the US economy (see, e.g., Medeiros et al., 2021; Goulet Coulombe et al., 2022; Clark et al., 2023; Goulet Coulombe, 2024), with datasets typically structured similarly to the one we use in this paper.
@@
Our approach to scenario analysis is developed under the assumption of a multivariate Gaussian reduced form error term with a general time-varying covariance matrix. From a practical and implementation perspective, we rely on a somewhat simplified version to capture time variation of the respective volatilities. The framework allows for flexible equation-by-equation estimation of the multivariate system and offers significant improvements in computational efficiency also in the context of CFs and GIRFs. In our empirical work, we capture heteroskedastic features of the data with a specification related to common volatility approaches (see Carriero et al., 2016), which reflects recent tools used to address outliers during the Covid-19 pandemic (see, e.g., Carriero et al., 2024). It is worth noting that all methods also work with more sophisticated volatility models (see, e.g., Chan, 2023), albeit at the cost of an increased computational burden.
@@
After briefly investigating the performance of our approach with synthetic data, we apply our framework in several related yet distinct cases. Our dataset comprises about $25$ quarterly macroeconomic and financial variables for the US economy ranging from the mid-$1970$s to the last quarter of $2023$. In one of our explorations, we also add some international variables and include data from the euro area (EA), Japan, and the United Kingdom (UK) in our model. The applied work assesses and illustrates the role of nonlinearities when interest centers on CFs, and we also explore asymmetries in the propagation of shocks of different types, signs and magnitudes. Specifically, we provide three empirical applications. First, inspired by Chan et al. (2025), we use a subset of the assumptions underlying the annual stress test conducted by the Federal Reserve System and compute CFs using soft constraints for different scenarios, comparing predictive densities from linear and nonlinear models. Second, reflecting the growth-at-risk literature (see Adrian et al., 2019), we study the unorthogonalized counterfactual implications of varying financial conditions (imposed via a hard constraint) on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock (as in, e.g., Barnichon et al., 2022), and compute GIRFs to shocks of different signs and magnitudes that are allowed to propagate internationally. We then use a restricted GIRF approach to gauge the role of spillovers and spillbacks.
@@
The rest of this paper is structured as follows. Section 2 discusses challenges and solutions for how to obtain predictive inference in the presence of nonlinearities of unknown form. We discuss how to impose constraints on forecasts, and how these constraints may be used to construct scenarios through GIRFs. Section 3 discusses an econometric implementation using BART. Section 4 contains three empirical illustrations. The last section concludes.
2. Nonlinearity and Predictive distributions
@@
Let $\bm{y}_{t}=(y_{1t},\ldots,y_{nt})^{\prime}$ collect $n$ variables for $t=1,\ldots,T,$ and $\bm{x}_{t}=(\bm{y}_{t-1}^{\prime},\ldots,\bm{y}_{t-p}^{\prime})^{\prime}$ is a $k=np$ vector of lags. Interest centers on multivariate models of the form: